Which of the following sequences would be represented by a linear function?

A.

4, 8, 12, 16...

B.

4, 16, 64, 256...

C.

1, 2, 4, 8...

D.

256, 128, 64, 32...
Answer #1

A sequence whose common difference is constant represents a linear function

in option A first term (a1) = 4;

second term (a2) =8

third term (a3) =12

fourth term (a4) = 16

common difference =(a4 -a3) = (a3-a2) = (a2-a1)

=(16-12) = (12-8) = (8-4) =4

so option A is correct

PLEASE LIKE IT RAISE YOUR THUMBS UP '

IF YOU ARE HAVING ANY DOUBT FEEL FREE TO ASK IN COMMENT SECTION